OpenStudy (anonymous):

Find the slope of a line that passes through the points (4, 5) and (4, –6). A. no slope B. 0 C. 11/8 D. -1/8

OpenStudy (anonymous):

Is B the correct answer?

OpenStudy (anonymous):

yes, its b

OpenStudy (anonymous):

or wait no

OpenStudy (anonymous):

Those two points have the same x coordinate, so you have a vertical line. A vertical line has no slope, A.

OpenStudy (anonymous):

what he said, the slope would result in division by 0

OpenStudy (nurali):

x1 =4 y1 =5 x2 = 4 y2 = -6 Slope, m = (y2 - y1) / (x2 - x1) m = (-6-5 / [4-4] m = -11/0 m = no slope